OpenStudy (anonymous):

Divide. (2x^2+9x+11)÷(x+2) Your answer should give the quotient and the remainder.

Nnesha (nnesha):

long division or synthetic which one is easy for you ?

OpenStudy (anonymous):

tbh I'm just not good at math at all @Nnesha

Nnesha (nnesha):

alright you will be an expert one day :-) alight bec leading coefficient is one of the divisor so let's do synthetic that one is easy :-) for that we have to solve x+2 =0 for x x+2 =0 x= ?

OpenStudy (anonymous):

um x=-2 ?

OpenStudy (sweetburger):



OpenStudy (sweetburger):

it literally tells u to divide by x+2

OpenStudy (sweetburger):

otherwise you would just use the rational root theorem

Nnesha (nnesha):

yeah.. 1)then carry down first term 2) multiply by -2 3)combine like terms done! at the end just put variables exponent of variable should be one less from the original question just like that^^^^

Nnesha (nnesha):

and that's how you should write 'n answer \[\huge\rm quotient + \frac{ remainder }{ divisor }\]